OpenStudy (anonymous):

[Calculus] Derivative of [tan(x)] / [1+cos(x)]?

OpenStudy (anonymous):

Quotient rule

OpenStudy (anonymous):

\[f(x) = \frac{ \tan(x) }{ 1+\cos(x) }\]

OpenStudy (anonymous):

\[\frac{ f'(x)g(x)-f(x)g'(x) }{ [g(x)]^2 }\]

OpenStudy (anonymous):

Start it off and tell us your answer, if you don't get it we will show you the steps where you messed up.